OpenStudy (anonymous):

a child's grandparents wish to purchase a bond that matures in 15 years.the bond pays an interest rate which is compounded semi-annually. if the bond's present value is $500,000 and it will be worth $1.2 million at maturity, a. what is the annual interest rate? b. if the bond company has a second option which offers an interest rate of 5% compounded continuously, which option should the grandparents choose?

OpenStudy (anonymous):

@hartnn ? :D

OpenStudy (anonymous):

@saifoo.khan ? @Hero ? @experimentX :D

OpenStudy (anonymous):

@.Sam. ? :D @estudier ? :D

OpenStudy (anonymous):

i'm confused on which m in A=P(1+r/m)^(mt) to use first?? since there's semi annually and annual?

OpenStudy (experimentx):

find the value of m first. Find the amount at the end of year and calculate the annual interest rate.

sam (.sam.):

compounded semi annually means money is compounded twice a year, so m = 2

sam (.sam.):

A=P(1+r/m)^(mt) Amount of money A which we will have at the end of t years if we invest an initial amount P at an annual rate r. \[1.2\times 10^6=(5\times 10^5)(1+\frac{r}{2})^{(2(15))}\] \[r \approx 0.0592245\]

OpenStudy (anonymous):

OHHHH annual interest rate is merely r.... okay silly moment right there.... hahaha thanks!